OpenStudy (anonymous):

Hey everyone, need help finding the arc length of the curve y^2=x^3 from (0,0) to (1/4,1/8).

OpenStudy (anonymous):

I know the formula and I think y'=(3x^2)/(2y)

OpenStudy (anonymous):

let me try this, see if it works. \[y=x^\frac{3}{2}\] yes?

OpenStudy (anonymous):

ok right, do we only take the +x^3/2?

OpenStudy (anonymous):

\[y'=\frac{3}{2}x^\frac{1}{2}\]

OpenStudy (anonymous):

y = x^(3/2) y' = 3/2*x^1/2 (y')^2 = 9/4*x integrate sqrt(1+9/4*x) from 1/8 to 1/4 use u-sub if you have to

OpenStudy (anonymous):

Rim would it not be from 0 to 1/4?

OpenStudy (anonymous):

oh sorry yes 0 to 1/4

OpenStudy (anonymous):

we are going from (0,0) to (1/8,1/4) so it is ok yes? maybe i am wrong, but i will graph it. in any case we have to integrate \[\int_0^{\frac{1}{8}} \sqrt{1-\frac{9}{4}x} dx\]

OpenStudy (anonymous):

oops you are right it is 1/4

OpenStudy (anonymous):

anti derivative is \[-\frac{(4-9x)^\frac{3}{2}}{27}\]

OpenStudy (anonymous):

and integral is \[\frac{8}{27}-\frac{7\sqrt{7}}{216}\]

OpenStudy (anonymous):

did you guys get .2824?

OpenStudy (anonymous):

let me get a calculator i will check

OpenStudy (anonymous):

no in fact i got .21055...

OpenStudy (anonymous):

ok, but isnt it also sqrt(1+(9/4)x)?

OpenStudy (anonymous):

yes it is .2824

OpenStudy (anonymous):

well i guess i messed up hold on

OpenStudy (anonymous):

you are right and i am wrong (0ften the case)

OpenStudy (anonymous):

take advantage of Wolfram alpha it is a very useful tool

OpenStudy (anonymous):

i put - where i should have had +. sorry

OpenStudy (anonymous):

no problem thanks for the help guys

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Just why do we pick positive sqrt all the time?